# Complements of caustics of the real $J_{10}$ singularities

### v3

@@ 2.4. The up-down involution
This action can also be extended to the corresponding virtual graphs. Namely, for any system of paths for the function $f(x,y)$ (see Fig. 1) we take the system of paths for f(x,y)-f(-x,y) obtained from it by the composition of the multiplication by 1-1 and the complex conjugation in ${\mathbb{C}}^{1}$. The virtual Morse function associated with the function f(x,y)-f(-x,y) and defined by the resulting system of paths is determined by that of the original virtual Morse function associated with $f(x,y)$, see § 2.2 in [16].
In particular, if two functions with ten real critical points are related by this involution, then their $D$-graphs are obtained from each other by replacing all minima by maxima and vice versa, and reversing the directions of all edges, cf. Proposition 16 in [16]
3. Enumeration of virtual components
Table 3. Numbers of virtual components of type $J_{10}^{1}$ (left) and $J_{10}^{3}$ (right)
@@ 3. Enumeration of virtual components
Table 3 shows the number of virtual components and the number of isotopy classes of Morse polynomials of type $\Phi_{1}$ and $\Phi_{3}$ with any value of the passport $(M,m_{+})$. The second of these numbers is given in parentheses when it differs from the first one. This table is detailed in Theorems 13, 14, 15.
@@
  • nine components of virtual Morse functions without local maxima; their D-graphs are shown in Figs. 59, and their Card invariants are indicated in the captions of these figures,
  • eleven virtual components obtained from the ones listed above by the involution ((5)) of the space $\Phi_{1}$.
@@ 2. There are exactly five virtual components of type $\Phi_{1}$ with eight real critical points:
  • one component with two local maxima and two minima, and Card=10890;
  • one component with exactly one local maximum, and Card=27378,
@@ 4. There are exactly three virtual components of type $\Phi_{1}$ with four real critical points:
5. There are exactly two virtual components of type $\Phi_{1}$ with two real critical points: one component without local maxima and one component without minima obtained from it by ((5)), both having Card=10608.
6. There is only one virtual component of type $\Phi_{1}$ without real critical points, its Card invariant is equal to 17642.
Figure 5. 97702 (left) and 93489 (right)
Figure 6. 68145 (left) and 42372 (right)
Figure 7. 52415 (left) and 63085 (right)
Figure 8. 82500 (left) and 27940 (right)
Figure 9. 33528
@@ 1. There are exactly 23 virtual components of type $\Phi_{3}$ with ten real critical points:
  • eleven virtual components obtained by the involution ((5)) from the components listed above ((except for the component with two minima and two maxima)).
@@ 2. There are exactly six virtual components of type $\Phi_{3}$ with eight real critical points:
  • one component with one local maximum, two minima and Card=66906,
@@
3. There are exactly three virtual components of type $\Phi_{3}$ with six real critical points:
  • one component with one local maximum, one minimum and Card=26922,
@@ 3. There are exactly three virtual components of type $\Phi_{3}$ with six real critical points:
4. There are exactly two virtual components of type $\Phi_{3}$ with four real critical points: one component without local maxima, having Card=21410, and one component obtained from the previous one by the involution ((5));
5. There is only one virtual component of type $\Phi_{3}$ with two real critical points: it has Card=14778 and no local extrema;
@@
Figure 13. 230472 (left) and 102234 (right)
Figure 14. 89320 (left) and 75108 (right)
Figure 15. 63756 (left) and 59862 (right)
Figure 16. 53130 (left) and 34870 (right)
Figure 17. 29370
@@
3.2. On the proofs of enumeration theorems
All possible values of the Card and D-graph invariants mentioned in Theorems 13 and 14 were found using the computer program described in [4], [13]. For any class $\Phi_{1}$ or $\Phi_{3}$, we use the Gusein-Zade–A’Campo method (see [1], [5]) to compute the intersection matrix of vanishing cycles of a Morse polynomial of this class having only real critical points. Then, using theorem 1.4 of §5.1 in [12], we calculate the intersection indices of these vanishing cycles with the set of real points. Thus, we obtain a virtual Morse function associated with this polynomial. Starting from this initial data, the combinatorial program
@@ 3.2. On the proofs of enumeration theorems
runs through the entire formal graph of this type and counts, in particular, the number of all virtual Morse functions of this type and the number of virtual functions with each passport invariant value. For each passport value for which this number is not zero, it then (upon request) finds a virtual Morse function with this value. Starting from this virtual Morse function, a slightly modified version of the same program (with virtual surgeries of types $s1$ and $s3$ disabled) counts the number of virtual Morse functions in its virtual component, i.e. the Card invariant of this virtual function. For example, this program
@@
with certain initial data handles the virtual component of the $\Phi_{3}$ class, which has eight real critical points and Card = 131148. If this number is smaller than the total number of virtual functions with that passport, we find another virtual function with the same passport but in a different virtual component. Then we calculate its Card value as well. We continue this process until the sum of the different values of the Card invariant of virtual functions with any given passport reaches the total number of virtual functions with that passport. This computation (and the subsequent reconstruction of D-graphs from the virtual Morse functions representing these components) proves all statements of Theorems 13 and 14.
In sections 5 and 6, we investigate the chirality of all these components and prove Theorem 15. Specifically, in § 5 we realize nearly all achiral virtual components by polynomials that are symmetric with respect to the reflection (4). In § 6.1 we describe a homological criterion of chirality in the terms of the formal graph and use it to demonstrate the achirality of the remaining achiral components. In § 6.2, we prove the chirality of all chiral components.
@@ 3.3. Normalization of $D$-graphs
a) it is directed from a vertex with a smaller Morse index to a vertex with a larger index,
b) it is dashed if the parities of Morse indices of corresponding critical points are the same, and is solid if these parities are different.
@@
b) seventeen out of eighteen $D$-graphs of polynomials without maxima shown in these figures split into pairs of standard Coxeter–Dynkin graphs of some simple singularities after normalization;
@@ 4. Proof of Proposition 9
The subset of the space $\Phi_{1}$ consisting of polynomials having a singularity of class $E_{8}$ with zero critical value, and also a singularity of class $A_{2}$, is the union of four connected components diffeomorphic to ${\mathbb{R}}^{7}$ and mapped to each other by the involutions ((4)) and ((5)). Two of these components consist of polynomials with negative critical values at the $A_{2}$ points, and the other two consist of polynomials with positive values. Each component is swept out by a one-parametric family of orbits of the group $\mathfrak{G}$}. The intersection of each such component with the parameter space of the deformation ((1)) is an orbit of the group $\mathbb{R}$$_{+}$ of positive quasihomogeneous dilations
@@
Proof. Let $f\in\Phi_{1}$ be a polynomial with a critical point of class $E_{8}$, in particular its 2-jet at this point is trivial and 3-jet has the form $q(x-\nu y)^{3},$ $q>0$. The substitution $\tilde{x}=x-\nu y$ (which belongs to the group $\mathfrak{G}$}) reduces it to the form $\tilde{x}^{3}$. The resulting function has zero coefficients at the monomials $y^{4}$ and $\tilde{x}y^{3}$, because otherwise its critical point 0 would be of class $E_{6}$ or $E_{7}$. Conversely, the coefficient at the monomial $y^{5}$ is non-zero: otherwise the Milnor number is greater than $8$. So, this function is a linear combination of monomials $x^{3},$ $x^{2}y^{2}$, $xy^{4}$, $y^{6}$, and $y^{5}$. An additional substitution of the form $\check{x}=\tilde{x}-\eta y^{2}$ (which also belongs to the group $\mathfrak{G}$}) reduces $f$ to the form
(7) $\check{x}^{3}+a\check{x}y^{4}\pm y^{6}+\tau y^{5}\quad\mbox{or}\quad\check{x}^{3}\pm\check{x}y^{4}+\tau y^{5},\qquad\tau\neq 0.$
The condition of having also a critical point of class $A_{2}$ (that is, the coincidence of two non-zero solutions of the system $f^{\prime}_{x}=0=f^{\prime}_{y}$) prohibits the second possibility and implies the condition $a=0$ in the first possibility of the alternative (7). The resulting set of polynomials of the form
(8) $x^{3}\pm y^{6}+\tau y^{5}$
splits into four connected components, which are characterized by the sign before the monomial $y^{6}$ and the sign of the coefficient $\tau$. These sets are all $\mathfrak{G}$}-invariant, and the critical values of their functions at the $A_{2}$ points are constant along $\mathfrak{G}$}-orbits. It is easy to show that the polynomials with different combinations of these signs cannot be $\mathfrak{G}$}-equivalent to each other. The group ${\mathbb{Z}}_{2}^{2}$ of two involutions (4), (5) maps all these four sets into each other. The orbits of the group $\mathfrak{G}$} that contain the functions (8) with different values of the parameter $t$ are distinct because they have different critical values at the points of class $A_{2}$. If a polynomial of the form (1) belongs to one of these sets, then its entire orbit under the action of the group (6) also belongs to it. $\Box$
@@
The subset of the space $\Phi_{3}$ consisting of polynomials having a singularity of class $D_{8}^{-}$ with zero critical value, and also a $A_{2}$ singularity, is the union of four connected components diffeomorphic to ${\mathbb{R}}^{7}$ and mapped to each other by the involutions ((4)) and ((5)). Two of these components consist of polynomials with the negative critical values at the $A_{2}$ points, and the other two consist of polynomials with the positive critical values. Each of these four components is swept out by a one-parametric family of orbits of the group $\mathfrak{G}$}. The intersection of each component with the parameter space of deformation ((2)) is an orbit of the group $\mathbb{R}$$_{+}$ of positive quasihomogeneous dilations ((6)).
Proof. Each function $f$ of class $\Phi_{3}$ with a critical point of class $D_{8}^{-}$ and critical value 0 can be reduced by the action of the group $\mathfrak{G}$} to the form
(9) $x^{3}+bx^{2}y^{2}\pm xy^{4}+\alpha x^{2}y.$
Indeed, the 2-jet of $f$ at this point is trivial, and the 3-jet has the form $q(x-\lambda y)^{2}(x-\nu y)$ for some $q>0$ and $\lambda\neq\nu$. The substitution $\tilde{x}=x-\lambda y$ and a dilation reduces this 3-jet to the form $\tilde{x}^{3}+\alpha\tilde{x}^{2}y$, $\alpha\neq 0$. The resulting function has zero coefficient at the monomial $y^{4}$: otherwise it would be of class $D_{5}$. The lower quasihomogeneous part with weights $\deg x=2,\deg y=1$ of the obtained polynomial should have the form $y(x-\theta y^{2})^{2}$, because otherwise the function is of the class $D_{6}$. So, this quasihomogeneous part can be reduced to the monomial $x^{2}y$ by a diffeomorphism of class $\mathfrak{G}$} (keeping the entire function in the class $\Phi_{3}$). The coefficient of the resulting polynomial at the monomial $y^{6}$ should vanish: otherwise the critical point is of class $D_{7}$. Finally, its coefficient at the monomial $xy^{4}$ is not zero, because otherwise the Milnor number would be greater than $8$. Dilating the coordinate $y$, we get the form (9).
The sum of Milnor numbers of all critical points of the obtained function is equal to ten, so there are only two critical points outside the origin (counting with the multiplicities). The system of equations $f^{\prime}_{x}=0=f^{\prime}_{y}$ for these two points implies easily the condition
@@
on their $y$-coordinate. If these two critical points coincide (at a point of class $A_{2}$), then $b^{2}\pm 21=0$, so the sign $\pm$ in (9) is equal to - and $b=\sqrt{21}$. The concluding arguments are the same as in the proof of Lemma 17. $\Box$
@@
The space of polynomials ((1)) contains exactly two polynomials having a critical point of class $E_{8}$ with zero critical value and a critical point of class $A_{2}$ with critical value 1. The space of polynomials ((2)) contains exactly two polynomials having a critical point of class $D_{8}^{-}$ with zero critical value and a critical point of class $A_{2}$ with critical value 1. In both cases, these two points are mapped to each other by the involution ((4)).

### v5

@@ 2.4. The up-down involution
This action can also be extended to the corresponding virtual Morse functions and D-graphs. Namely, for any system of paths for the function $f(x,y)$ (see Fig. 1) we take the system of paths for f(x,y)-f(-x,y) obtained from it by the composition of the multiplication by 1-1 and the complex conjugation in ${\mathbb{C}}^{1}$. The virtual Morse function associated with the function f(x,y)-f(-x,y) and defined by the resulting system of paths is determined by that of the original virtual Morse function associated with $f(x,y)$, see § 2.2 in [21].
In particular, if two functions with ten real critical points are related via this involution, then their $D$-graphs are obtained from each other by replacing all minima by maxima and vice versa, and reversing the directions of all edges, see Proposition 16 in [21].
3. Enumeration of isotopy classes and virtual components
Table 3. Numbers of virtual and real components of class $\Phi_{1}$ (left) and $\Phi_{3}$ (right)
@@ 3. Enumeration of isotopy classes and virtual components
Table 3 shows the number of virtual components and the number of isotopy classes of Morse polynomials of types $\Phi_{1}$ and $\Phi_{3}$ with any value of the passport $(M,m_{+})$. The second number is given in parentheses when it differs from the first. This table is detailed in Theorems 14, 15, 16.
@@
There are exactly 2005366 different virtual Morse functions of $J_{10}^{1}$ class distributed as follows into 37 virtual components.
@@
  • nine components of virtual Morse functions without local maxima: their D-graphs are shown in Figs. 59, and their Card invariants are indicated in the captions of these figures, and
  • eleven virtual components obtained from the ones listed above in this theorem via the involution ((5)) of the space $\Phi_{1}$.
@@ 2. There are exactly five virtual components of type $\Phi_{1}$ with eight real critical points:
  • one component with two local maxima, two minima, and Card=10890;
  • one component with exactly one local maximum and Card=27378,
@@ 4. There are exactly three virtual components of type $\Phi_{1}$ with four real critical points:
  • a virtual component with no local minima obtained from the previous component via involution ((5)).
5. There are exactly two virtual components of type $\Phi_{1}$ with two real critical points: one component without local maxima and one component without local minima obtained from it by involution ((5)), both having Card=10608.
6. There is only one virtual component of class $\Phi_{1}$ without real critical points, its Card invariant is equal to 17642.
Figure 5. $\Phi_{1}$, no maxima; 97702 (left) and 93489 (right)
Figure 6. $\Phi_{1}$, no maxima; 68145 (left) and 42372 (right)
Figure 7. $\Phi_{1}$, no maxima; 52415 (left) and 63085 (right)
Figure 8. $\Phi_{1}$, no maxima; 82500 (left) and 27940 (right)
Figure 9. $\Phi_{1}$, no maxima; 33528
@@
There are exactly 2970134 different virtual Morse functions of $J_{10}^{3}$ type distributed as follows into 35 virtual components.
@@ 1. There are exactly 23 virtual components of type $\Phi_{3}$ with ten real critical points:
  • eleven virtual components obtained by the involution ((5)) from all the components listed above in this theorem except for the component with two minima and two maxima.
@@ 2. There are exactly six virtual components of type $\Phi_{3}$ with eight real critical points:
  • one component with one local maximum, two local minima, and Card=66906,
@@
3. There are exactly three virtual components of type $\Phi_{3}$ with only six real critical points:
  • one component with one local maximum, one local minimum, and Card=26922,
@@ 3. There are exactly three virtual components of type $\Phi_{3}$ with only six real critical points:
4. There are exactly two virtual components of type $\Phi_{3}$ with only four real critical points: one component without local maxima, having Card=21410, and one component obtained from it by the involution ((5));
5. There is only one virtual component of type $\Phi_{3}$ with exactly two real critical points: it has Card=14778 and no local extrema;
@@
Figure 13. $\Phi_{3}$, no maxima; 230472 (left) and 102234 (right)
Figure 14. $\Phi_{3}$, no maxima; 89320 (left) and 75108 (right)
Figure 15. $\Phi_{3}$, no maxima; 63756 (left) and 59862 (right)
Figure 16. $\Phi_{3}$, no maxima; 53130 (left) and 34870 (right)
Figure 17. $\Phi_{3}$, no maxima; 29370
@@
3.2. On the proofs of enumeration theorems 14 and 15
All possible values of the Card and D-graph invariants mentioned in Theorems 14 and 15 were found using the computer program described in [5], [18]. For any type $\Phi_{1}$ or $\Phi_{3}$, we use the Gusein-Zade–A’Campo method (see [1], [6]) to compute the intersection matrix of vanishing cycles of a Morse polynomial of this class having only real critical points. Then, using theorem 1.4 of §5.1 in [17], we calculate the intersection indices of these vanishing cycles with the set of real points. Thus, we obtain a virtual Morse function associated with this polynomial. Starting from this initial data, the combinatorial program
@@ 3.2. On the proofs of enumeration theorems 14 and 15
runs through the entire formal graph of this type and counts, in particular, the number of all virtual Morse functions of this type and the number of virtual functions with each passport invariant value. For each passport value for which this number is not zero, it then (upon request) finds a virtual Morse function with this value. Starting from this virtual Morse function, a slightly modified version of the same program (with virtual surgeries of types $s1$ and $s3$ disabled) counts the number of virtual Morse functions in its virtual component, i.e., the Card invariant of this virtual function. (For example, this program
@@
with certain initial data handles the virtual component of the $\Phi_{3}$ class, which has eight real critical points and Card = 131148.) If this number is smaller than the total number of virtual functions with that passport, we find another virtual function with the same passport but in a different virtual component. Then we calculate its Card value as well. We continue this process until the sum of the different values of the Card invariant of virtual functions with any given passport reaches the total number of virtual functions with that passport. This computation (and the subsequent reconstruction of D-graphs from the virtual Morse functions representing these components in the case of polynomials with ten real critical points) proves all statements of Theorems 14 and 15.
In sections 5 and 6, we investigate the chirality of all these components and prove Theorem 16. Specifically, in § 5 we realize nearly all achiral virtual components by polynomials that are symmetric with respect to the reflection (4). In § 6.1, we describe a homological criterion of chirality in the terms of the formal graph and use it to demonstrate the achirality of the remaining achiral components. In § 6.2, we prove the chirality of all chiral components.
@@ 3.3. Normalization of $D$-graphs
a) it is oriented from a vertex with a smaller Morse index to a vertex with a larger index,
b) it is dashed if the parities of Morse indices of the critical points corresponding to its ends are the same; it is solid if these parities are different.
@@
b) seventeen out of eighteen $D$-graphs of polynomials without maxima shown in these figures are split by the normalization into pairs of standard Coxeter–Dynkin graphs of some simple singularities after normalization;
@@ 4. Proof of Proposition 9
The subset of the space $\Phi_{1}$, which consists of polynomials having a critical point of class $E_{8}$ with zero critical value, and also a critical point of class $A_{2}$, is the union of four smooth components diffeomorphic to ${\mathbb{R}}^{7}$ and moved to each other by the involutions ((4)) and ((5)). Two of these components consist of polynomials with negative critical values at the $A_{2}$ points, and the other two consist of polynomials with positive values. Each component is swept out by a one-parametric family of orbits of the group $\mathfrak{G}$}. Each of these orbits contains exactly one polynomial of the form
(6) $x^{3}\pm y^{6}+\tau y^{5},\quad\tau\neq 0\ .$
The intersection of each such component with the parameter space of the deformation ((1)) is an orbit of the group $\mathbb{R}$$_{+}$ of positive quasihomogeneous dilations
@@
Proof. Let $f$ be a polynomial of type $\Phi_{1}$ having a real critical point of class $E_{8}$ with zero critical value. Its 2-jet at this point is trivial, and 3-jet has the form $q(x-\nu y)^{3},$ $q>0$. The substitution $\tilde{x}=x-\nu y$ and a dilation of $x$ coordinate reduce this 3-jet to the form $\tilde{x}^{3}$. The resulting polynomial $\tilde{f}\in\Phi_{1}$ has zero coefficients at the monomials $y^{4}$ and $\tilde{x}y^{3}$, because otherwise its critical point at the origin would be of class $E_{6}$ or $E_{7}$. Conversely, the coefficient at the monomial $y^{5}$ is non-zero: otherwise the Milnor number is greater than $8$. So, this polynomial is a linear combination of monomials $\tilde{x}^{3},$ $\tilde{x}^{2}y^{2}$, $\tilde{x}y^{4}$, $y^{6}$, and $y^{5}$. An additional substitution of the form $\check{x}=\tilde{x}-\eta y^{2}$ and dilation of $y$ coordinate (which belong to the group $\mathfrak{G}$}) reduce $\tilde{f}$ to the form
(8) $\check{x}^{3}+a\check{x}\check{y}^{4}\pm\check{y}^{6}+\tau\check{y}^{5}\qquad\mbox{or}\qquad\check{x}^{3}\pm\check{x}\check{y}^{4}+\tau\check{y}^{5},$
$\tau\neq 0.$ The condition of having also a critical point of class $A_{2}$ (that is, the coincidence of two non-zero complex solutions of the system $f^{\prime}_{x}=0=f^{\prime}_{y}$) prohibits the right-hand possibility of the alternative (8) and implies the condition $a=0$ in the left-hand possibility. The resulting set of polynomials of the form (6) consists of four components, which are characterized by the sign before the monomial $y^{6}$ and the sign of the coefficient $\tau$.
All points (6) belong to different $\mathfrak{G}$}-orbits. Indeed, the critical value at the $A_{2}$-point is an invariant of orbits. This value is a monomial of degree six of the parameter $\tau$ in (6). In particular, for any $c\neq 0$ there are exactly two polynomials (6) having the critical value equal to $c$. These two two polynomials are obtained from each other by the multiplication of $\tau$ by 1-1. They are not $\mathfrak{G}$}-equivalent because the $y$ coordinates of their $A_{2}$ points have different signs. So, the subset of $\Phi_{1}$ considered in Lemma is the family of $\mathfrak{G}$}-orbits represented and parametrized by polynomials (6). The group ${\mathbb{Z}}_{2}^{2}$ of two involutions (4), (5) maps all its four components into each other. If a polynomial of the form (1) belongs to one of these components, then its entire orbit under the action of the one-parametric group (7) also belongs to it. These one-dimensional orbits are smooth curves in the space of polynomials (1), and the $\mathfrak{G}$}-orbits intersect this space transversally. Therefore, the four components in consideration are canonically diffeomorphic to the products of these smooth curves and the group $\mathfrak{G}$}. $\Box$
@@
The subset of the space $\Phi_{3}$, which consists of polynomials having a critical point of class $D_{8}^{-}$ with zero critical value, and also a class $A_{2}$ critical point, is the union of four connected components that are diffeomorphic to ${\mathbb{R}}^{7}$ and are moved to each other by the involutions ((4)) and ((5)). Two of these components consist of polynomials with the negative critical values at the $A_{2}$ critical points, and the other two consist of polynomials with the positive critical values. Each of these four components is swept out by a one-parametric family of orbits of the group $\mathfrak{G}$}. Each of these orbits contains exactly one polynomial of the form
(9) $x^{3}\pm\sqrt{21}x^{2}y^{2}-xy^{4}+\alpha x^{2}y\ .$
The intersection of each component with the parameter space of deformation ((2)) is an orbit of the group $\mathbb{R}$$_{+}$ of positive quasihomogeneous dilations ((7)).
Proof. Each polynomial $f$ of type $\Phi_{3}$ with a critical point of class $D_{8}^{-}$ and critical value 0 at this point can be reduced by the action of the group $\mathfrak{G}$} to the form
(10) $x^{3}+bx^{2}y^{2}\pm xy^{4}+\alpha x^{2}y,\quad\alpha\neq 0.$
Indeed, the 2-jet of $f$ at its $D_{8}^{-}$ critical point is trivial, and the 3-jet has the form $q(x-\varkappa y)^{2}(x-\nu y)$ for some $q>0$ and $\varkappa\neq\nu$. The substitution $\tilde{x}=x-\varkappa y$ and a dilation of $\tilde{x}$ coordinate reduce this 3-jet to the form $\tilde{x}^{3}+\alpha\tilde{x}^{2}y$, $\alpha\neq 0$. The resulting polynomial has zero coefficient at the monomial $y^{4}$: otherwise this polynomial would be of class $D_{5}$. Its lower quasihomogeneous part with weights $\deg\tilde{x}=2,\deg y=1$ should have the form $y(\tilde{x}-\theta y^{2})^{2}$, because otherwise the polynomial is of the class $D_{6}$. So, this quasihomogeneous part can be reduced to the monomial $\tilde{x}^{2}y$ by a diffeomorphism of class $\mathfrak{G}$} keeping the entire polynomial within the space $\Phi_{3}$. The coefficient of the resulting polynomial at the monomial $y^{6}$ should vanish, because otherwise the critical point is of class $D_{7}$. Finally, the coefficient at the monomial $\tilde{x}y^{4}$ is not zero, because otherwise the Milnor number would be greater than $8$. Dilating the coordinate $y$, we get a polynomial of the form (10).
The sum of Milnor numbers of all critical points in ${\mathbb{C}}^{2}$ of the obtained polynomial is equal to ten, so there are only two critical points outside the origin (counting with the multiplicities). The system of equations $f^{\prime}_{x}=0=f^{\prime}_{y}$ for these two points implies easily the condition
@@
on their $y$-coordinate, where the sign at $48y^{2}$ repeats the sign at $xy^{4}$ in (10). If these two critical points coincide (at a point of class $A_{2}$), then $b^{2}\pm 21=0$, so the sign $\pm$ in (10) is equal to - and $b$ is equal to $\sqrt{21}$ or $\sqrt{21}$. The critical values of the resulting polynomials (9) at the $A_{2}$ point depend on the parameter $\alpha$ as a monomial of degree six. They are negative in the case of the sign ++ at $\sqrt{21}x^{2}y^{2}$ and are positive in the case of the sign -. The concluding arguments are the same as in the proof of Lemma 18. $\Box$
@@
1. The space of polynomials ((1)) contains exactly two polynomials having a critical point of class $E_{8}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{1}{5}$. Namely, these two polynomials are the intersection points of this space with the $\mathfrak{G}$}-orbits of the polynomials $x^{3}+y^{6}\pm\frac{6}{5}y^{5}$.
2. The space of polynomials ((2)) contains exactly two polynomials having a critical point of class $D_{8}^{-}$ with zero critical value and a critical point of class $A_{2}$ with critical value $\frac{343343}{375000}\sqrt{\frac{7}{3}}$. Namely, these two polynomials are the intersection points of this space and the $\mathfrak{G}$}-orbits of polynomials $x^{3}+\sqrt{21}x^{2}y^{2}-xy^{4}\pm 6x^{2}y.$
In both cases, the two polynomials are mapped to each other by the involution ((4)). $\Box$
Proof. Proof follows immediately from Lemmas 18 and 19, Proposition 1, and direct calculations with polynomials (6) and (9).
Figure 18. Morse perturbations of $E_{8}$ and $D_{8}^{-}$ singularities
Lemma 21.
Each of the two polynomials considered in statement 1 of Corollary 20 has arbitrarily small Morse perturbations which
a)) split its critical point of class $E_{8}$ into four saddlepoints with zero critical value and four minima in such a way that the local zero-level set of the obtained polynomial looks as shown in Fig. 18 ((left)) or its reflection about a vertical line, and
b)) split the $A_{2}$ critical point into one minimum point and one saddlepoint with critical values near $\frac{1}{5}$.
The $D$-graph of these Morse perturbations is shown in Fig. 6 ((left)).
Lemma 22.
Each of the two polynomials considered in statement 2 of Corollary 20 has arbitrarily small Morse perturbations which
a)) split its critical point of class $D_{8}^{-}$ into five saddlepoints with zero critical value and three minima in such a way that the local zero-level set of the obtained polynomial looks as shown in Fig. 18 ((right)) or its reflection with respect to a vertical line, and
b)) split the $A_{2}$ critical point into one minimum point and one saddlepoint with critical values near $\frac{343343}{375000}\sqrt{\frac{7}{3}}$.
The $D$-graph of these Morse perturbations is shown in Fig. 15 ((right)).
Proofs of Lemmas 21, 22. The existence of such independent perturbations of two critical points follows from the versality of the deformations (1), (2) and from the perturbations of $E_{8}$ and $D_{8}^{-}$ singularities demonstrated in pp. 17 and 15 of [1]. By the construction, the vertices of the D-graph of the first (respectively, the second) perturbation can be split into subsets of cardinality eight and two in such a way that the corresponding subgraphs are the canonical Coxeter-Dynkin graphs of types $E_{8}$ (respectively, $D_{8}$) and $A_{2}$. Among all D-graphs listed in Theorem 14 (respectively, 15), only the D-graph called in the last assertion of Lemma 21 (respectively, Lemma 22) allows such a splitting. $\Box$
The proof of Proposition 9 follows from the statements 1822 in the same way as Proposition 3 of [20] is deduced from Lemmas 4 and 5 and Proposition 22. First, we describe this proof for the type $\Phi_{1}$ polynomials, referring to [20] for the common details.